OpenStudy (anonymous):

Sam bought brushes for $8, a palette for $5, and oil paints for $15. He paid $29.82 in all. What sales-tax rate did Sam pay?

OpenStudy (kamille):

8+5+15=28$ 28$<===>x% 29,82$<====>100% x=(28*100)/29.82=93, 90 sales - tax = 100%-93,90=6,10%